OpenStudy (anonymous):

It's early in the evening, and you're standing in your back yard waiting for the stars to come out. Before long, you can see two stars in the sky. The equation of the line between you and Venus (which is actually a planet, but looks like a star) is y = 10x + 30. The equation of the line between you and Sirius is y = 9x + 40. Which star is higher in the sky? Why?

OpenStudy (anonymous):

@$w3G_Godd

OpenStudy (anonymous):

Hey man, this is the last one for the night.

OpenStudy (anonymous):

Sorry about all the questions...

OpenStudy (anonymous):

its np one sec

OpenStudy (anonymous):

OK. :)

OpenStudy (anonymous):

y = 10x + 30 y = 9x + 40 2/2y = 19x + 70 / 2 y=19x+35 are you doing linear equations? because if you are you take the y and multiply it by the x and then add the 35, and then plot it on the graph paper or w/e you're doing it on. thats how i learned to do it, this year.

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

Sirius, the y intercept is greater, suggesting that the "planet" is higher up in the sky.

OpenStudy (anonymous):

And correct, Liner Equations.

OpenStudy (anonymous):

ok :)